OpenStudy (anonymous):

Your daily workout plan involves a total of 60 minutes of running and swimming. You burn 20 calories per minute running and 10 calories per minute swimming. How many calories will you burn in your 60−minute workout if you run for 30 minutes?

OpenStudy (anonymous):

30*20=600 30*10=300 600+300=900

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem.